# A characterization of zero entropy loosely Bernoulli flows via FK-pseudometric

### v1

@@ 1. Introduction
Despite the strong connection between continuous and discrete-time dynamics, it is sometimes not immediate to find direct analogies between these settings. For example, there is a notion of expansivity, whose definition for homeomorphisms turns out to be unfit for continuous flows. Indeed, no non-trivial flow satisfies the direct translation of the discrete-time definition to flows [BW72]. Therefore a lot of work has been done to define expansivity for flows so that it behaves similarly to the discrete-time notion. In this work we propose a definition of a pseudometric for flows corresponding to the Feldman-Katok pseudometric introduced in [KŁ17] for transformations (discrete-time systems).
Two topologically conjugated discrete-time systems have the same dynamics from the topological perspective, Unfortunately, in the continuous-time case topological conjugacy seems to be too restrictive. A very small time-change of a continuous flow might not be conjugated to the initial one. In this setting a weaker form of equivalence called (topological) orbit equivalence seems to be more natural. Orbit equivalence guarantees similarity between a flow and its time-changes. However, it fails to preserve certain topological properties and quantities such as mixing and entropy.
The equivalence relations can also be considered from the measure-theoretical viewpoint and equivalent systems are supposed to share the same dynamics from the perspective of ergodic theory. Unfortunately, the measurable orbit equivalence becomes too weak. Dye, in [Dye59] and [Dye63], showed that any two ergodic flows (or automorphisms) are measurably orbit equivalent.
@@
In the opposite direction, a remarkable result on the isomorphism problem, due to Ornstein (see [Orn70a], [Orn70b], [Orn70c]) says that Bernoulli shifts with the same entropy are isomorphic. In 1976, studying the isomorphism problem in ergodic theory Feldman [Fel76] introduced the definition of loosely Bernoulli automorphisms inspired by Ornstein’s definition of very weak Bernoulli from [Orn70b]. In both cases, one first defines the property for finite partitions and then for automorphisms of measure space. While the notion of very weak Bernoulli automorphism is based on the normalized Hamming distance $\bar{d}_{n}$ between strings of length $n$, the notion of loose Bernoullicity for automorphisms is based on the weaker edit metric $\bar{f}_{n}$. These distances are related to dynamical pseudometrics for discrete-time systems, the first one called Besicovitch pseudometric and the second one the Feldman-Katok pseudometric (see [KŁO17], [KŁ17] for details).
@@
Roughly speaking, Kakutani [Kak43] conjectured that all zero-entropy systems were loosely Kronecker. The conjecture turned out to be false and the first example of a zero-entropy but not loosely Kronecker system was given by Feldman [Fel76] followed by results of Katok [Kat77] and Ornstein, Rudolph and Weiss [ORW82] who independently constructed uncountably many zero entropy systems that are not Kakutani equivalent one to another
@@
In the early 80’s there emerged results dealing directly with flows. The first one is due to Feldman [Fel80] who proved among many other things a characterization of loosely Kronecker flows. Inspired by his work, Ratner [Rat81] introduced invariants of Kakutani equivalence, that were used to show that different cartesian powers of horocycle flows are not Kakutani equivalent.
In her last paper [Rat17], Ratner presented a slightly different definition of her invariants from [Rat81]. Inspired by Ratner’s definition which can be seen as a continuous-time version of the $\bar{f_{n}}$-metric we define the Feldman-Katok pseudometric (FK-pseudometric for short) for flows.
@@
Let $\Phi$ be a continuous flow on $X$ and $\mu$ be an ergodic $\Phi$-invariant measure. The measure-preserving flow $(X,\Phi,\mu)$ is loosely Kronecker if and only if there exists a Borel set $H\subset X$ such that $\mu(H)=1$ and $\operatorname{\tilde{\rho}_{FK}}(x,y)=0$ for every $x,y\in H$.
@@
For both, discrete and continuous-time case, the connection between measure-preserving and continuous (topological) systems is well established. On one hand, the Krylov-Bogoliubov Theorem guarantees the existence of an invariant measure for continuous flows and maps. On the other hand, Jewett-Krieger Theorem says that every ergodic transformation is isomorphic to a uniquely ergodic continuous map. We call a continuous flow a topological model of a measure-preserving flow if it is uniquely ergodic and isomorphic to the measure-preserving one with respect to its unique ergodic invariant measure. For flows, Denker and Eberlein [DE74] obtained a result similar to the Jewett-Krieger Theorem proving that every ergodic measure-preserving flow admits a minimal topological model.
@@
Let $\Phi$ be a continuous flow and $\operatorname{{\mathcal{M}}_{\Phi}^{e}}(X)$ be the set of all ergodic $\Phi$-invariant measures, then the following conditions are equivalent:
  1. (1)
    $\Phi$ is topologically loosely Kronecker;
  2. (2)
    $\Phi$ is uniquely ergodic and $(X,\Phi,\mu)$ is loosely Kronecker, where $\{\mu\}~{}=~{}\operatorname{{\mathcal{M}}_{\Phi}^{e}}(X)$.
@@
Organization of the paper:
  • Section 2: We present the basic definitions about flows and ergodic theory that appear in this work.
  • Section 3: We recall the definition of the FK-pseudometric for the discrete case and its properties, then we introduce our definition for flows stipulating parallels between the two.
  • Section 4: We recall Ratner’s characterization of loosely Kronecker flows and prove Theorem 1.
  • Section 5: We present the notion of topologically loosely Kronecker flows and prove the Theorem 2.
  • Section 6: We provide examples of topologically loosely Kronecker flows and we present some consequences of Theorem 1 and Theorem 2.
@@ 2. Basic definitions
Throughout this work by $X$ we mean a compact metric space and $d$ is a compatible metric. Whenever we mention measure-theoretical properties of $X$ we will consider $X$ to be a measurable space endowed with its Borel $\sigma$-algebra. The Lebesgue measure on the space of real numbers $\mathbb{R}$ will be denoted by $\lambda$. The set of positive real numbers will be denoted by $\mathbb{R}^{+}$.
@@ 2.1. Flows
A flow $\Phi$ on $X$ is a map $\Phi\colon\mathbb{R}\times X\to X$ such that $\Phi(0,x)=x$ and $\Phi(t+s,x)=\Phi(s,\Phi(t,x))$ for all $x\in X$ and $t,s\in\mathbb{R}$. If the map $\Phi$ is continuous with respect to the product topology on ×$\mathbb{R}\times X$ we say it is a continuous flow. For $t\in\mathbb{R}$ the time-t map is defined as $\varphi^{t}:=\Phi(t,\cdot)\colon X\to X$. We call $\Phi$ a measurable flow if $\varphi^{t}$ is measurable for every $t\in\mathbb{R}$. Since $\varphi^{0}$ is simply the identity, we will tacitly assume $t\neq 0$ when referring to time-t maps.
The forward orbit of $x$ under $\Phi$ is ${\mathcal{O}}^{+}(x):=\left\{\varphi^{t}(x)\in X\mid t\geq 0\right\}$ and the backward orbit of $x$ under $\Phi$ is ${\mathcal{O}}^{-}(x):=\left\{\varphi^{t}(x)\in X\mid t\leq 0\right\}$. The orbit of $x$ under $\Phi$ is ${\mathcal{O}}(x):={\mathcal{O}}^{+}(x)\cup{\mathcal{O}}^{-}(x)$.
A singularity of a flow is a point $x\in X$ such that ${\mathcal{O}}(x)=\{x\}$. In this work we will only consider flows without singularities.
A continuous flow is:
  • transitive if for some $x\in X$, the set ${\mathcal{O}}^{+}(x)$ is dense in $X$,
  • minimal if for every $x\in X$, the set ${\mathcal{O}}^{+}(x)$ is dense in $X$.
Let $T\colon X\to X$ be a homeomorphism, $\alpha\colon X\to\mathbb{R}^{+}$ a continuous function and the map $\tau:X\times\mathbb{R}\to X\times\mathbb{R}$ given by $\tau(x,s)=(T(x),s-\alpha(x))$. We denote by $X^{\alpha}_{T}=X\times\mathbb{R}\big{/}\sim$ the quotient space where $(x,s)\sim\tau^{n}(x,s)$ for every $n\in\mathbb{Z}$.
The special flow over the transformation $T$ with roof function $\alpha$ is the flow on $X^{\alpha}_{T}$ induced by the time translation $f^{t}(x,s)=(x,s+t)$. In case $\alpha\equiv 1$ we call it the suspension flow over $T$ and we denote $X^{1}_{T}=X_{T}$. It is possible to define a metric on $X_{T}$ induced by $d$ assuming without loss of generality that $(X,d)$ has diameter at most 1 (see [Tho82] for details). For $x,y\in X$ and $t\in[0,1)$ such metric is given by:
@@ 2.2. Ergodic Theory
By ${\mathcal{M}}(X)$ we denote the set of all Borel probability measures on $X$. Let $\Phi$ be a measurable flow on $X$. We say that a measure $\mu\in{\mathcal{M}}(X)$ is $\Phi$-invariant if it is $\varphi^{t}$-invariant for every $t\in\mathbb{R}$. The set of all $\Phi$-invariant measures will be denoted by $\operatorname{{\mathcal{M}}_{\Phi}}(X)$.
We say that an invariant measure $\mu\in\operatorname{{\mathcal{M}}_{\Phi}}(X)$ is ergodic if for any measurable subset $A\subset X$ with $\varphi^{t}(A)=A$ for all $t\in\mathbb{R}$ either $\mu(A)=0$ or $\mu(A)=1$. The set of ergodic measures will be denoted by $\operatorname{{\mathcal{M}}_{\Phi}^{e}}(X)$.
It is well known that ${\mathcal{M}}(X)$ endowed with the so called weak* topology is a compact metrizable space and $\operatorname{{\mathcal{M}}_{\Phi}}(X)\subset{\mathcal{M}}(X)$ a compact subspace. It is also well known that such topology is induced by the Prokhorov metric given by:
$D_{P}(\mu,\nu)=\inf\{\varepsilon>0:\mu(B)\leq\nu(B^{\varepsilon})+\varepsilon\text{ for every Borel set }B\subset X\},$
where $B^{\varepsilon}$ denotes the $\varepsilon$-hull of $B$, that is $B^{\varepsilon}=\{y\in X:d(y,B)<\varepsilon\}$.
Let $\delta_{x}$ denote the Dirac measure supported on $\{x\}$ for $x\in X$. For $t\in\mathbb{R}^{+}$ the t-empirical measure of $x\in X$ (with respect to $\Phi$) will be denoted by:
@@
Two measure-preserving flows $(X,\Phi,\mu)$ and $(Y,\Psi,\nu)$ are isomorphic if there exists a measure-preserving isomorphism $h\colon X\to Y$ such that $h\circ\varphi^{t}=\psi^{t}\circ h$ for every $t\in\mathbb{R}$ and $\mu$-almost every $x\in X$. It is worth mentioning Ambrose-Kakutani Representation Theorem which says that every measure-preserving flow is isomorphic to a special flow.
@@
We say that a flow $\Psi$ on $X$ is a time-change of $\Phi$ if there exists $\alpha\in L^{1}_{+}(X,\mu)$ such that:
@@
We say that two measure-preserving flows are Kakutani equivalent (also called monotonically equivalent) if one of them is isomorphic to a time-change of the other.
@@
A measure-preserving flow $(X,\Phi,\mu)$ is called loosely Kronecker if it is Kakutani equivalent to Kronecker flow.
@@ 3.1. Discrete Feldman-Katok pseudometric
Let us recall the pseudometric with respect to a continuous map $T\colon X\to X$. Fix $x,y\in X$, $\delta>0$ and $n\in\mathbb{N}$. We define an $(n,\delta)$-matching between $x$ and $y$ to be an order preserving bijection $\pi\colon D\to D^{\prime}$ such that $D,D^{\prime}\subset\{0,...,n-1\}$ and for every $i\in D$ we have $d(T^{i}(x),T^{\pi(i)}(y))<\delta$. Given an $(n,\delta)$-matching $\pi\colon D\to D^{\prime}$, let $|\pi|$ be the cardinality of $D$.
@@
The $(n,\delta)$-gap between $x$ and $y$ is given by:
@@
The discrete-time $f_{\delta}$-pseudometric is given by:
@@
The discrete-time Feldman-Katok pseudometric is given by:

### v2

@@ 1. Introduction
Despite the strong connection between continuous and discrete-time dynamics, it is sometimes not immediate to find direct analogies between these settings. For example, the definition of expansivity for homeomorphisms turns out to be unfit for continuous flows. Indeed, no non-trivial flow satisfies the direct translation of the discrete-time definition [BW72]. Therefore, a lot of work has been done to define expansivity for flows so that it behaves similarly to the discrete-time notion. The first definition of expansivity that appeared for flows (due to Bowen and Walters [BW72]), for example, preserves certain results known for homeomorphisms, but it does not work well with flows with singularities, where the so-called $k^{*}$-expansivity (due to Komuro [Kom84]) is more suitable. In summary, in order to find an analogue definition, several proposals were made, and each one of them preserves different results known for homeomorphisms on different classes of flows (see Section 1.7 in [FH19]). In this work we propose a definition of a pseudometric for flows corresponding to the Feldman-Katok pseudometric introduced in [KŁ17] for discrete-time systems and we show that it preserves results obtained in [GRK22].
Two topologically conjugated discrete-time systems have the same dynamics from the topological perspective. Unfortunately, in the continuous-time case topological conjugacy seems to be too restrictive. A very small time-change of a continuous flow might not be conjugated to the initial one. In this setting a weaker form of equivalence called (topological) orbit equivalence seems to be more natural. Orbit equivalence guarantees similarity between a flow and its time-changes. However, it fails to preserve certain topological properties and quantities such as mixing and entropy. The equivalence relations can also be considered from the measure-theoretical viewpoint and equivalent systems are supposed to share the same dynamics from the perspective of ergodic theory. Unfortunately, the measurable orbit equivalence becomes too weak. Dye, in [Dye59] and [Dye63], showed that any two ergodic automorphisms are measurably orbit equivalent and his result can be extended to the context of ergodic flows by using cross-sections.
@@
In the opposite direction, a remarkable result on the isomorphism problem, due to Ornstein (see [Orn70a], [Orn70b], [Orn70c]) says that Bernoulli shifts with the same entropy are isomorphic. In 1976, studying the isomorphism problem in ergodic theory Feldman [Fel76] introduced the definition of loosely Bernoulli automorphisms inspired by Ornstein’s definition of very weak Bernoulli from [Orn70b]. One first defines the property for finite partitions and then for automorphisms. While the notion of very weak Bernoulli partition is based on the normalized Hamming distance $\bar{d}_{n}$ between strings of length $n$, the notion of loose Bernoullicity is based on the weaker edit metric $\bar{f}_{n}$. These distances are related to dynamical pseudometrics for discrete-time systems, the first one called Besicovitch pseudometric and the second one the Feldman-Katok pseudometric (see [KŁO17], [KŁ17] for details).
@@
Roughly speaking, Kakutani [Kak43] conjectured that all zero entropy systems were loosely Kronecker. The conjecture turned out to be false and the first example of a zero entropy but not loosely Kronecker system was given by Feldman [Fel76] followed by results of Katok [Kat77] and Ornstein, Rudolph and Weiss [ORW82] who independently constructed uncountably many zero entropy systems that are not Kakutani equivalent one to another.
@@
In the early 80’s there emerged results dealing directly with flows. The first one is due to Feldman [Fel80] who proved among many other things a characterization of loosely Kronecker flows. Inspired by his work, Ratner [Rat81] introduced invariants of Kakutani equivalence that were used to show that different cartesian powers of horocycle flows are not Kakutani equivalent.
In her last paper [Rat17], Ratner presented a slightly different definition of her invariants from [Rat81]. Inspired by Ratner’s definition which can be seen as a continuous-time version of the $\bar{f_{n}}$-metric, we define the Feldman-Katok pseudometric (FK-pseudometric for short) for flows.
@@
Let $\Phi$ be a continuous flow on $X$ and $\mu$ be an ergodic $\Phi$-invariant measure. The measure-preserving flow $(X,\Phi,\mu)$ is loosely Kronecker if and only if there exists a Borel set $H\subseteq X$ such that $\mu(H)=1$ and $\operatorname{\tilde{\rho}_{FK}}(x,y)=0$ for every $x,y\in H$.
@@
For both, discrete and continuous-time case, the connection between measure-preserving and continuous (topological) systems is well established. On one hand, the Krylov-Bogoliubov Theorem guarantees the existence of an invariant measure for continuous flows and maps. On the other hand, the Jewett-Krieger Theorem says that every ergodic transformation is isomorphic to a uniquely ergodic continuous map. We call a continuous flow a topological model of a measure-preserving flow if it is uniquely ergodic and isomorphic to the measure-preserving one with respect to its unique ergodic invariant measure. For flows, Denker and Eberlein [DE74] obtained a result similar to the Jewett-Krieger Theorem proving that every ergodic measure-preserving flow admits a minimal topological model.
@@
Let $\Phi$ be a continuous flow and $\operatorname{{\mathcal{M}}_{\Phi}^{e}}(X)$ be the set of all ergodic $\Phi$-invariant measures. Then $\Phi$ is topologically loosely Kronecker if and only if $\Phi$ is uniquely ergodic and $(X,\Phi,\mu)$ is loosely Kronecker, where $\{\mu\}~{}=~{}\operatorname{{\mathcal{M}}_{\Phi}^{e}}(X)$.
@@
Here is the organization of the paper. In section 2 we present the basic definitions about flows and ergodic theory that appear in this work. In section 3 we recall the definition of the FK-pseudometric for the discrete case and its properties, then we introduce our definition for flows stipulating parallels between the two. In section 4 we recall Ratner’s characterization of loosely Kronecker flows and prove Theorem 1. In section 5 we present the notion of topologically loosely Kronecker flows and prove the Theorem 2. In section 6 we provide examples of topologically loosely Kronecker flows and we present some consequences of Theorem 1 and Theorem 2.
@@ 2. Basic definitions
Throughout this work by $X$ we mean a compact metrizable space and $d$ is a compatible metric. Whenever we mention measure-theoretical properties of $X$ we will consider $X$ to be a measurable space endowed with its Borel $\sigma$-algebra. The Lebesgue measure on the space of real numbers $\mathbb{R}$ will be denoted by $\lambda$. The set of non-negative real numbers will be denoted by $\mathbb{R}^{+}$.
@@ 2.1. Flows
A flow $\Phi$ on $X$ is a map $\Phi\colon\mathbb{R}\times X\to X$ such that $\Phi(0,x)=x$ and $\Phi(t+s,x)=\Phi(s,\Phi(t,x))$ for all $x\in X$ and $t,s\in\mathbb{R}$. If the map $\Phi$ is continuous (resp. measurable) with respect to the product topology (resp. product $\sigma$-algebra) on ×$\mathbb{R}\times X$ we say it is a continuous flow (resp. measurable flow). For $t\in\mathbb{R}$ the time-$t$ map is defined as $\varphi^{t}:=\Phi(t,\cdot)\colon X\to X$. Since $\varphi^{0}$ is simply the identity, we will tacitly assume $t\neq 0$ when referring to time-$t$ maps.
The forward orbit of $x$ under $\Phi$ is ${\mathcal{O}}^{+}(x):=\left\{\varphi^{t}(x)\in X:t\in\mathbb{R}^{+}\right\}$ and the backward orbit of $x$ under $\Phi$ is ${\mathcal{O}}^{-}(x):=\left\{\varphi^{-t}(x)\in X:t\in\mathbb{R}^{+}\right\}$. The orbit of $x$ under $\Phi$ is ${\mathcal{O}}(x):={\mathcal{O}}^{+}(x)\cup{\mathcal{O}}^{-}(x)$. A continuous flow is transitive (resp. minimal) if for some (resp. every) $x\in X$, the set ${\mathcal{O}}^{+}(x)$ is dense in $X$. A singularity of a flow is a point $x\in X$ such that ${\mathcal{O}}(x)=\{x\}$. In this work we will only consider flows without singularities.
Let $T\colon X\to X$ be a homeomorphism, $\alpha\colon X\to\mathbb{R}^{+}\backslash\{0\}$ a continuous function and the map $\tau:X\times\mathbb{R}\to X\times\mathbb{R}$ given by $\tau(x,s)=(T(x),s-\alpha(x))$. We denote by $X^{\alpha}_{T}=X\times\mathbb{R}\big{/}\sim$ the quotient space where $(x,s)\sim\tau^{n}(x,s)$ for every $n\in\mathbb{Z}$.
The special flow over the transformation $T$ with roof function $\alpha$ is the flow on $X^{\alpha}_{T}$ induced by the time translation $f^{t}(x,s)=(x,s+t)$. In case $\alpha\equiv 1$ we call it the suspension flow over $T$ and we denote $X^{1}_{T}=X_{T}$. It is possible to define a metric $d_{T}$ on $X_{T}$ induced by $d$ assuming without loss of generality that $(X,d)$ has diameter at most 1 (see [Tho82] for details). For $t\in[0,1)$, the restriction of $d_{T}$ to $X\times\{t\}$ is given for $x,y\in X$ by
@@ 2.2. Ergodic Theory
By ${\mathcal{M}}(X)$ we denote the set of all Borel probability measures on $X$. The support of a measure $\mu\in{\mathcal{M}}(X)$ is $\operatorname{supp}(\mu):=\{x\in X:\mu(U)>0\text{ for all $U\subseteq X$ open neighborhood of }x\}$. Let $\Phi$ be a measurable flow on $X$. We say that a measure $\mu\in{\mathcal{M}}(X)$ is $\Phi$-invariant if it is $\varphi^{t}$-invariant for every $t\in\mathbb{R}$. The set of all $\Phi$-invariant measures will be denoted by $\operatorname{{\mathcal{M}}_{\Phi}}(X)$.
We say that an invariant measure $\mu\in\operatorname{{\mathcal{M}}_{\Phi}}(X)$ is ergodic if for any measurable subset $A\subseteq X$ with $\varphi^{t}(A)=A$ for all $t\in\mathbb{R}$ either $\mu(A)=0$ or $\mu(A)=1$. The set of ergodic measures will be denoted by $\operatorname{{\mathcal{M}}_{\Phi}^{e}}(X)$.
It is well known that ${\mathcal{M}}(X)$ endowed with the so-called weak* topology is a compact metrizable space and $\operatorname{{\mathcal{M}}_{\Phi}}(X)\subseteq{\mathcal{M}}(X)$ a compact subspace. It is also well known that such topology is induced by the Prokhorov metric given by
$D_{P}(\mu,\nu)=\inf\{\varepsilon>0:\mu(B)\leq\nu(B^{\varepsilon})+\varepsilon\text{ for every Borel set }B\subseteq X\},$
where $B^{\varepsilon}$ denotes the $\varepsilon$-hull of $B$, that is, $B^{\varepsilon}=\{y\in X:d(y,B)<\varepsilon\}$.
Let $\delta_{x}$ denote the Dirac measure supported on $\{x\}$ for $x\in X$. For $t\in\mathbb{R}^{+}$ the $t$-empirical measure of $x\in X$ (with respect to $\Phi$) will be denoted by
@@
Definition 2.1.
We say that two measure-preserving flows $(X,\Phi,\mu)$ and $(Y,\Psi,\nu)$ are isomorphic if there exists a measure-preserving isomorphism $h\colon X\to Y$ such that $(h\circ\varphi^{t})(x)=(\psi^{t}\circ h)(x)$ for every $t\in\mathbb{R}$ and $\mu$-almost every $x\in X$.
It is worth mentioning the Ambrose-Kakutani Representation Theorem which says that every measure-preserving flow is isomorphic to a special flow.
@@
We say that a flow $\Psi$ on $X$ is a time-change of $\Phi$ if there exists $\alpha\in L^{1}_{+}(X,\mu)$ such that
@@
We say that two measure-preserving flows are Kakutani equivalent if one of them is isomorphic to a time-change of the other.
@@
A measure-preserving flow $(X,\Phi,\mu)$ is called loosely Kronecker if it is Kakutani equivalent to a Kronecker flow.
@@ 3.1. Discrete Feldman-Katok pseudometric
Let us recall the pseudometric with respect to a continuous map $T\colon X\to X$. Fix $x,y\in X$, $\delta>0$ and $n\in\mathbb{N}$. We define an $(n,\delta)$-matching between $x$ and $y$ to be an order preserving bijection $\pi\colon D\to D^{\prime}$ such that $D,D^{\prime}\subseteq\{0,\ldots,n-1\}$ and for every $i\in D$ we have $d(T^{i}(x),T^{\pi(i)}(y))<\delta$. Given an $(n,\delta)$-matching $\pi\colon D\to D^{\prime}$, let $|\pi|$ be the cardinality of $D$.
@@
The $(n,\delta)$-gap between $x$ and $y$ is given by
@@
The discrete-time $f_{\delta}$-pseudometric is given by
@@
The discrete-time Feldman-Katok pseudometric is given by